OpenStudy (anonymous):

April adds the expressions 9x + 7 and 5x – 3. Evan adds the expressions 14x + 2 and 4x + 3. How much greater is Evan’s sum than April’s sum? x +

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

Well lets add April's terms 9x+7+5x-3 =

myininaya (myininaya):

collect like terms... 9x+5x=? 7-3?

OpenStudy (anonymous):

4

myininaya (myininaya):

7-3=4 and 9x+5x=?

OpenStudy (anonymous):

14x

myininaya (myininaya):

very good so april's sum is 14x+4 now lets look at evan's

myininaya (myininaya):

14x+2+4x+3=?

myininaya (myininaya):